ELECTRONIC TRANSITIONS AND SPIN-ORBIT SPLITTING OF LANTHANUM DIMER

\underline{YANG LIU}, LU WU, CHANGHUA ZHANG, and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Lanthanum dimer (La_{2}) was studied by mass-analyzed threshold ionization (MATI) spectroscopy and a series of highlevel multi-configuration ab initio calculations (CASSCF, CASPT2, and MRCI). The MATI spectrum exhibits three band systems originating at 39044, 40312, and 40862 cm^{−1}, respectively. Above the band origin, the first band system displays a vibrational progression of ∼232 cm^{−1}, and the other two show vibrational progression with the same interval of ∼240 cm^{−1}. Below the band origin, the three systems exhibit the same vibrational interval of ∼207 cm^{−1}. These band systems are assigned to three electronic transitions from the ground state of La_{2} to the low-lying electronic states of La_{2}^{+}: ^{2}Σ_{g}^{+} ← ^{1}Σ_{g}^{+}, ^{2}Π_{u,1/2} ← ^{1}Σ_{g}^{+}, and ^{2}Π_{u,3/2} ← ^{1}Σ_{g}^{+}. The spin-orbit splitting in the ^{2}Π_{u} ion state is 550 cm^{−1}. In addition, the electronic states and bonding of La_{2} will be compared with those of Sc_{2} and Y_{2}.